# Matching conditions at null infinity in the presence of logarithms: the role of advanced and retarded radiation

### v1

@@ 1 Introduction
The purpose of this paper is to provide new insight on the matching conditions through a different, physically motivated explanation for the emergence of logarithmic terms in the asymptotic expansion at null infinity of massless scalar and electromagnetic fields in Minkowski space. Our approach is complementary to that of Fuentealba:2024lll ; Fuentealba:2025ekj in that it relies from the very begining on the behaviour of the fields near null infinity rather than on their Cauchy development from initial data.
@@
The connection between advanced waves and non-power-law radial expansions of the fields near $\mathscr{I}^{+}$ violating the peeling property of $\mathscr{I}^{+}$ has been recognized since the earliest studies of gravitational radiation at null infinity. In fact, the standard power-law expansion of the leading orders in the asymptotic expansion in the BMS formalism might be considered to be equivalent to Sommerfeld’s radiation condition, which is required to describe radiation emitted by bounded sources (see, e.g., Bondi:1962px ; Sachs:1961zz ; Sachs:1962wk ; Tamburino:1966zz ; Kroon:1998dv ). For this reason, such terms were typically neglected due to the imposition of the Sommerfeld radiation condition. Nonetheless, there could be physically relevant scenarios where incoming radiation, and consequently leading logarithmic terms, could play a crucial role. This occurs for instance in the study of scattering processes Kraus:2025wgi .
We go beyond these earlier observations on the connection between non-polynomial terms at null infinity and advanced radiation by making the explicit connection with the initial data on Cauchy hyperplanes and the resulting Cauchy development, obtaining thereby a new physical insight on the generalized matching conditions valid when these logarithmic terms, involving both $\log(-u)$ and $\log r$, are present.
@@ 4 Conclusions
Remarkably, this advanced/retarded wave approach enables the determination of the matching conditions for the radial logarithmic terms between the future of past null infinity and the past of future null infinity. This is achieved by evaluating the fields in the vicinity of $i^{0}$ and imposing that their asymptotic behaviour be of the Coulombic type, free of $\log r$ contributions at spatial infinity. The matching conditions derived using this approach are in exact agreement with those in Refs. Fuentealba:2024lll ; Fuentealba:2025ekj .
@@
It is natural to ask to what extent these results can be extended to the gravitational case. For linearized gravity, one may naturally expect a behaviour closely analogous to that of the scalar and electromagnetic fields discussed in this work. However, in the fully nonlinear theory, the situation is considerably more subtle. This is because, as observed for the scalar and electromagnetic fields discussed above, all subleading terms in the radial expansion near $\mathscr{I}^{-}$ contribute to the leading $\log r$ term at $\mathscr{I}^{+}$. In General Relativity, it is extremely challenging to control all subleading terms in the radial expansion due to the inherent nonlinearities of the theory. However, since Bondi’s energy-loss formula Bondi:1960jsa has a similar form to that of the scalar and electromagnetic cases, the requirement of finite total energy implies that the shear tensor at late advanced times must take the form $C_{AB}\underset{v\rightarrow\infty}{=}C_{AB}^{+}\log v+\dots$. Consequently, when it is reexpressed in retarded coordinates, it will generically give rise to $\log r$ contributions in the asymptotic expansion near $\mathscr{I}^{+}$. How to connect these logarithmic terms to the polyhomogeneous solutions involving dominant logarithmic terms presented for example, in Refs. Winic:1985 ; Chrusciel:1993hx ; Andersson:1994ng remains an interesting question for future investigation.

### v2

@@ 1 Introduction
The purpose of this paper is to provide new insight on the matching conditions through a different, physically motivated explanation for the emergence of logarithmic terms in the asymptotic expansion at null infinity of massless scalar and electromagnetic fields in Minkowski space. Our approach is complementary to that of Fuentealba:2024lll ; Fuentealba:2025ekj in that it relies from the very beginning on the behaviour of the fields near null infinity rather than on their Cauchy development from initial data.
@@
The connection between advanced waves and non-power-law radial expansions of the fields near $\mathscr{I}^{+}$ violating the peeling property of $\mathscr{I}^{+}$ has been recognized since the earliest studies of gravitational radiation at null infinity. In fact, the standard power-law expansion of the leading orders in the asymptotic expansion in the BMS formalism might be considered to be equivalent to Sommerfeld’s radiation condition, which is required to describe radiation emitted by bounded sources (see, e.g., Bondi:1962px ; Sachs:1961zz ; Sachs:1962wk ; Tamburino:1966zz ; Kroon:1998dv ). For this reason, such terms were typically neglected due to the imposition of the Sommerfeld radiation condition. Nonetheless, there could be physically relevant scenarios where incoming radiation saturating the flux bound, and consequently leading logarithmic terms, could play a crucial role. This occurs for instance in the study of the $S$-matrix where Arefeva-Faddeev-Slavnov type of boundary conditions Arefeva:1974jv ; Faddeev:1980be are relevant Kim:2023qbl ; Kraus:2025wgi . In particular, a $\log r$ dominant term in the angular component of the vector potential involving the Feynman propagator has been explicitly exhibited in Kim:2023qbl (appendix D, formula (D.8) and sentence below).
Recent articles Sahoo:2018lxl ; Campiglia:2019wxe ; AtulBhatkar:2019vcb ; AtulBhatkar:2021txo ; Compere:2025tzr have also addressed the problem of matching conditions in the presence of logarithms, confirming the “mixed matching conditions phenomenon" obtained by integrating the fields from initial data given on Cauchy hypersurfaces. However, these interesting papers do not treat the leading logarithms that motivate the present work. Of course, one recovers the conditions studied in these papers as a special case if one sets the coefficients of the leading logarithms equal to zero.
@@ 4 Conclusions
Remarkably, this advanced/retarded wave approach enables the determination of the matching conditions for the radial logarithmic terms between the future of past null infinity and the past of future null infinity. This is achieved by evaluating the fields in the vicinity of $i^{0}$ and imposing that their asymptotic behaviour be of the Coulombic type, free of $\log r$ contributions at spatial infinity. The matching conditions derived using this approach are in exact agreement with those in Refs. Fuentealba:2024lll ; Fuentealba:2025ekj . Our derivation not only confirms, but also provides a new physical insight on these matching conditions.
@@
One might wonder whether the dominant logarithmic terms found at null infinity do not spoil convergence of the symplectic form and of the angular momentum. While we have not directly studied this question at null infinity, the issue was indirectly addressed on Cauchy hypersurfaces (spatial infinity) in Henneaux:2018gfi ; Henneaux:2018hdj ; Henneaux:2018mgn , where it was shown that no problem arises if only the dominant $P$-branch or the dominant $Q$-branch is present. Indeed both branches decay with exactly the same powers of $r$ at spatial infinity, and the $Q$-branch does not exhibit $\log r$ term. Each branch is separately manifestly consistent on Cauchy hypersurfaces. It is only when both are simultaneously present that divergences appear [^6] In this respect, it is somewhat intriguing, and in any case consistent with finiteness, that the (imaginary part of the) solution of Kim:2023qbl , appendix D, has a pure dominant $Q$-term.. Since the ADM angular momentum at spatial infinity is finite for a pure dominant $Q$ branch, it provides a finite initial condition at null infinity ($u\rightarrow$-$\infty$) for the angular momentum even in that case, guaranteeing a finite flux as long as the angular momentum remains finite at later $u$’s. It would be of interest to fully explore this question.
It is natural to ask to what extent the results of this paper can be extended to the gravitational case. For linearized gravity, one may naturally expect a behaviour closely analogous to that of the scalar and electromagnetic fields discussed in this work. However, in the fully nonlinear theory, the situation is considerably more subtle. This is because, as observed for the scalar and electromagnetic fields discussed above, all subleading terms in the radial expansion near $\mathscr{I}^{-}$ contribute to the leading $\log r$ term at $\mathscr{I}^{+}$. In General Relativity, it is extremely challenging to control all subleading terms in the radial expansion due to the inherent nonlinearities of the theory. However, since Bondi’s energy-loss formula Bondi:1960jsa has a similar form to that of the scalar and electromagnetic cases, the requirement of finite total energy implies that the shear tensor at late advanced times must take the form $C_{AB}\underset{v\rightarrow\infty}{=}C_{AB}^{+}\log v+\dots$. Consequently, when it is reexpressed in retarded coordinates, it will generically give rise to $\log r$ contributions in the asymptotic expansion near $\mathscr{I}^{+}$. How to connect these logarithmic terms to the polyhomogeneous solutions involving dominant logarithmic terms presented for example, in Refs. Winic:1985 ; Chrusciel:1993hx ; Andersson:1994ng remains an interesting question for future investigation.